It took Fran 5.4 to drive to her mothers house on saturday morning. On sunday it took her 6 hours to get home. Her average speed was 7 mph slower, what was the average speed on sunday?
Answer by lwsshak3(3711)

It took Fran 5.4 to drive to her mothers house on saturday morning. On sunday it took her 6 hours to get home. Her average speed was 7 mph slower, what was the average speed on sunday
**
let x=Fran's average speed on Sunday
x+7=Fran's average speed on Saturday
distance=travel time*speed (same for both days)
..
5.4(x+7)=6x
5.4x+37.8=6x
.6x=37.8
x=37.8/.6=63
ans:
Fran's average speed on Sunday=63 mph